# Científicas (CSIC) and Universitat de València (UV), 46980 Valencia, Spain

### v1

@@ B.1 Warsaw Basis
[…]
$\mathcal{C}_{\rm{i}}$ $\Lambda^{\rm{eff}}_{\rm{i}}$ on-pole Obs. on-pole $\Lambda^{\rm{eff}}_{\rm{i}}$ off-pole Obs. off-pole $\Lambda^{\rm{eff}}_{\rm{i}}$ Combined
$\mathcal{C}_{HWB}$ 111.55 $A^{Z}_{e}$ 5.88 $\sigma_{Zh}^{240}$ 111.55
$[\mathcal{C}_{Hl}^{(3)}]_{22}$ 67.88 $A^{Z}_{e}$ 22.09 $\sigma_{WW}^{240}$ 68.06
$[\mathcal{C}_{Hl}^{(3)}]_{11}$ 65.12 $A^{Z}_{\mu}$ 23.76 $\sigma_{WW}^{240}$ 65.39
$\mathcal{C}_{HD}$ 57.11 $m_{W}$ 3.69 $\sigma_{WW}^{240}$ 57.11
$[\mathcal{C}_{ll}]_{1221}$ 52.98 $A^{Z}_{e}$ 31.80 $R_{\mu}^{163}$ 54.66
$[\mathcal{C}_{He}]_{11}$ 54.23 $A^{Z}_{e}$ 11.38 $\sigma_{Zh}^{365}$ 54.25
$[\mathcal{C}_{He}]_{22}$ 53.58 $A^{Z}_{\mu}$ 1.17 $R_{\mu}^{240}$ 53.58
$[\mathcal{C}_{Hl}^{(1)}]_{22}$ 51.18 $A^{Z}_{\mu}$ 0.79 $R_{\mu}^{240}$ 51.18
$[\mathcal{C}_{Hl}^{(1)}]_{11}$ 50.85 $A^{Z}_{e}$ 12.63 $\sigma_{Zh}^{365}$ 50.89
$[\mathcal{C}_{lq}^{(1)}]_{1133}$ 12.54 $A^{Z}_{e}$ 41.92 $R_{b}^{240}$ 41.79
$[\mathcal{C}_{lq}^{(1)}]_{1122}$ 2.03 $A^{Z}_{e}$ 41.00 $R_{c}^{240}$ 41.09
$[\mathcal{C}_{lq}^{(3)}]_{1111}$ 6.84 $A^{Z}_{e}$ 40.88 $R_{c}^{240}$ 40.99
$[\mathcal{C}_{lq}^{(3)}]_{1133}$ 17.14 $A^{Z}_{\mu}$ 40.83 $R_{b}^{240}$ 40.84
$[\mathcal{C}_{lq}^{(3)}]_{1122}$ 6.81 $A^{Z}_{e}$ 39.10 $R_{c}^{240}$ 39.44
$[\mathcal{C}_{He}]_{33}$ 37.27 $A^{Z}_{\tau}$ 1.17 $R_{\tau}^{240}$ 37.27
$[\mathcal{C}_{Hl}^{(3)}]_{33}$ 36.59 $A^{Z}_{\tau}$ 1.82 $R_{\tau}^{163}$ 36.59
$[\mathcal{C}_{Hl}^{(1)}]_{33}$ 35.72 $A^{Z}_{\tau}$ 0.79 $R_{\tau}^{240}$ 35.72
$[\mathcal{C}_{eu}]_{1111}$ 3.26 $A^{Z}_{e}$ 31.53 $R_{c}^{240}$ 31.62
$[\mathcal{C}_{Hq}^{(3)}]_{11}$ 31.01 $\Gamma_{Z}$ 1.68 $\sigma_{Zh}^{365}$ 31.01
$[\mathcal{C}_{Hq}^{(1)}]_{33}$ 30.81 $m_{W}$ 1.58 $\sigma_{WW}^{240}$ 30.81
$[\mathcal{C}_{ll}]_{1122}$ 11.94 $A^{Z}_{e}$ 29.75 $R_{\mu}^{163}$ 30.22
$[\mathcal{C}_{ll}]_{1133}$ 2.30 $A^{Z}_{e}$ 29.75 $R_{\tau}^{163}$ 30.03
$[\mathcal{C}_{ll}]_{1331}$ 2.25 $A^{Z}_{e}$ 29.66 $R_{\tau}^{163}$ 29.89
$[\mathcal{C}_{dH}]_{33}$ 0.00 29.82 $\mu_{b\bar{b}}$ 29.82
$[\mathcal{C}_{ee}]_{1122}$ 2.67 $A^{Z}_{e}$ 27.92 $R_{\mu}^{240}$ 27.73
$[\mathcal{C}_{ee}]_{1133}$ 2.38 $A^{Z}_{e}$ 27.92 $R_{\tau}^{240}$ 27.73
$[\mathcal{C}_{ed}]_{1133}$ 2.29 $A^{Z}_{e}$ 26.37 $R_{b}^{240}$ 26.32
$[\mathcal{C}_{Hu}]_{33}$ 26.04 $m_{W}$ 1.63 $\sigma_{WW}^{240}$ 26.04
$[\mathcal{C}_{Hq}^{(3)}]_{22}$ 25.60 $R^{Z}_{\mu}$ 1.64 $\sigma_{Zh}^{365}$ 25.60
$[\mathcal{C}_{eu}]_{1122}$ 3.26 $A^{Z}_{e}$ 25.45 $R_{e}^{365}$ 25.54
$[\mathcal{C}_{eH}]_{33}$ 0.00 24.39 $\mu_{\tau\bar{\tau}}$ 24.39
$[\mathcal{C}_{Hq}^{(3)}]_{33}$ 23.70 $R^{Z}_{\mu}$ 1.49 $R_{b}^{240}$ 23.70
$[\mathcal{C}_{qe}]_{3311}$ 13.22 $A^{Z}_{e}$ 22.97 $R_{b}^{163}$ 23.54
$[\mathcal{C}_{uW}]_{33}$ 22.30 $A^{Z}_{e}$ 1.06 $\sigma_{Zh}^{365}$ 22.30
$[\mathcal{C}_{uB}]_{33}$ 22.05 $A^{Z}_{e}$ 0.88 $\sigma_{Zh}^{365}$ 22.06
$[\mathcal{C}_{eH}]_{22}$ 0.00 21.77 $\mu_{\mu\bar{\mu}}$ 21.77
$[\mathcal{C}_{lu}]_{1111}$ 3.04 $A^{Z}_{e}$ 20.48 $R_{c}^{240}$ 20.41
$[\mathcal{C}_{uH}]_{22}$ 0.00 19.89 $\mu_{c\bar{c}}$ 19.89
$[\mathcal{C}_{eu}]_{1133}$ 12.39 $A^{Z}_{e}$ 18.90 $R_{t}^{365}$ 19.69
$[\mathcal{C}_{lu}]_{1133}$ 11.78 $A^{Z}_{e}$ 18.07 $R_{t}^{365}$ 18.80
$[\mathcal{C}_{le}]_{2211}$ 2.65 $A^{Z}_{e}$ 18.17 $R_{\mu}^{240}$ 18.16
$[\mathcal{C}_{le}]_{3311}$ 2.38 $A^{Z}_{e}$ 18.17 $R_{\tau}^{240}$ 18.15
$[\mathcal{C}_{le}]_{1122}$ 2.62 $A^{Z}_{\mu}$ 18.15 $R_{\mu}^{240}$ 18.13
$[\mathcal{C}_{le}]_{1133}$ 2.29 $A^{Z}_{e}$ 18.15 $R_{\tau}^{240}$ 18.13
$[\mathcal{C}_{ed}]_{1122}$ 2.28 $A^{Z}_{e}$ 17.94 $R_{e}^{365}$ 17.92
$[\mathcal{C}_{qe}]_{1111}$ 2.30 $A^{Z}_{e}$ 17.89 $R_{e}^{365}$ 17.88
$[\mathcal{C}_{ed}]_{1111}$ 2.29 $A^{Z}_{e}$ 17.87 $R_{e}^{365}$ 17.85
$[\mathcal{C}_{lq}^{(3)}]_{2233}$ 17.57 $A^{Z}_{e}$ 5.69 $\sigma_{WW}^{240}$ 17.71
$[\mathcal{C}_{Hq}^{(1)}]_{11}$ 17.64 $\Gamma_{Z}$ 0.60 $\sigma_{Zh}^{365}$ 17.64
$[\mathcal{C}_{ld}]_{1133}$ 2.16 $A^{Z}_{e}$ 17.42 $R_{b}^{365}$ 17.43
$[\mathcal{C}_{lu}]_{1122}$ 3.04 $A^{Z}_{e}$ 16.68 $R_{e}^{365}$ 16.65
$[\mathcal{C}_{qe}]_{2211}$ 2.29 $A^{Z}_{e}$ 16.41 $R_{e}^{365}$ 16.41
$[\mathcal{C}_{ll}]_{1111}$ 1.06 $\Gamma_{Z}$ 16.53 $R_{e}^{365}$ 16.33
$[\mathcal{C}_{ee}]_{1111}$ 4.39 $A^{Z}_{e}$ 16.37 $R_{e}^{365}$ 16.22
$[\mathcal{C}_{Hq}^{(1)}]_{22}$ 15.06 $\Gamma_{Z}$ 0.84 $R_{c}^{240}$ 15.06
$[\mathcal{C}_{le}]_{1111}$ 3.12 $A^{Z}_{e}$ 13.59 $R_{e}^{365}$ 13.59
$[\mathcal{C}_{Hu}]_{11}$ 13.17 $R^{Z}_{\mu}$ 1.00 $R_{c}^{240}$ 13.17
$[\mathcal{C}_{Hu}]_{22}$ 13.10 $R^{Z}_{\mu}$ 0.84 $\sigma_{Zh}^{365}$ 13.10
$[\mathcal{C}_{lq}^{(1)}]_{1111}$ 2.04 $A^{Z}_{e}$ 12.58 $R_{e}^{365}$ 13.09
$[\mathcal{C}_{qe}]_{3322}$ 13.07 $A^{Z}_{\mu}$ 1.71 $R_{\mu}^{240}$ 13.07
$[\mathcal{C}_{lq}^{(1)}]_{2233}$ 12.61 $A^{Z}_{\mu}$ 1.22 $R_{t}^{365}$ 12.61
$[\mathcal{C}_{eu}]_{2233}$ 12.26 $A^{Z}_{\mu}$ 2.34 $R_{\mu}^{240}$ 12.27
$[\mathcal{C}_{ld}]_{1122}$ 2.16 $A^{Z}_{e}$ 11.88 $R_{e}^{365}$ 11.89
$[\mathcal{C}_{ld}]_{1111}$ 2.16 $A^{Z}_{e}$ 11.83 $R_{e}^{365}$ 11.85
$[\mathcal{C}_{lu}]_{2233}$ 11.81 $A^{Z}_{\mu}$ 1.58 $R_{\mu}^{240}$ 11.82
$[\mathcal{C}_{quqd}^{(1)}]_{3333}$ 0.00 11.79 $\mu_{b\bar{b}}$ 11.79
$[\mathcal{C}_{Hd}]_{11}$ 10.85 $\Gamma_{Z}$ 0.59 $\sigma_{Zh}^{365}$ 10.85
$[\mathcal{C}_{dW}]_{33}$ 1.38 $A^{Z}_{e}$ 8.93 $\mu_{b\bar{b}}$ 10.35
$\mathcal{C}_{HW}$ 9.29 $A^{Z}_{e}$ 7.76 $\sigma_{Zh}^{365}$ 10.22
$[\mathcal{C}_{Hd}]_{33}$ 9.64 $R^{Z}_{\mu}$ 0.83 $R_{b}^{240}$ 9.65
$[\mathcal{C}_{eH}]_{11}$ 0.00 9.63 $\mu_{\mu\bar{\mu}}$ 9.63
$\mathcal{C}_{HB}$ 9.16 $A^{Z}_{e}$ 3.95 $\sigma_{Zh}^{240}$ 9.23
$[\mathcal{C}_{qe}]_{3333}$ 9.10 $A^{Z}_{\tau}$ 1.71 $R_{\tau}^{240}$ 9.10
$[\mathcal{C}_{lq}^{(1)}]_{3333}$ 8.84 $A^{Z}_{\tau}$ 1.22 $R_{t}^{365}$ 8.84
$[\mathcal{C}_{lequ}^{(1)}]_{3333}$ 0.00 8.82 $\mu_{\tau\bar{\tau}}$ 8.82
$\mathcal{C}_{W}$ 8.63 $A^{Z}_{e}$ 4.61 $\sigma_{WW}^{240}$ 8.79
$[\mathcal{C}_{qq}^{(1)}]_{3333}$ 8.71 $R^{Z}_{\mu}$ 1.87 $R_{t}^{365}$ 8.72
$[\mathcal{C}_{eu}]_{3333}$ 8.53 $A^{Z}_{\tau}$ 2.34 $R_{\tau}^{240}$ 8.57
$[\mathcal{C}_{Hd}]_{22}$ 8.48 $R^{Z}_{\mu}$ 0.60 $\sigma_{Zh}^{365}$ 8.48
$[\mathcal{C}_{lu}]_{3333}$ 8.25 $A^{Z}_{\tau}$ 1.58 $R_{\tau}^{240}$ 8.25
$[\mathcal{C}_{lq}^{(3)}]_{3333}$ 7.65 $A^{Z}_{\tau}$ 4.53 $R_{\tau}^{163}$ 8.25
$[\mathcal{C}_{lequ}^{(1)}]_{2233}$ 0.00 7.87 $\mu_{\mu\bar{\mu}}$ 7.87
$[\mathcal{C}_{lq}^{(3)}]_{2211}$ 6.82 $A^{Z}_{\mu}$ 4.65 $R_{\mu}^{240}$ 7.70
$[\mathcal{C}_{lq}^{(3)}]_{2222}$ 6.79 $A^{Z}_{\mu}$ 4.64 $R_{\mu}^{240}$ 7.67
$\mathcal{C}_{H\square}$ 6.84 $m_{W}$ 3.30 $\sigma_{Zh}^{240}$ 6.92
$[\mathcal{C}_{qq}^{(3)}]_{1111}$ 5.42 $\Gamma_{Z}$ 4.59 $R_{c}^{240}$ 6.79
$[\mathcal{C}_{qu}^{(8)}]_{2332}$ 0.00 6.78 $\mu_{c\bar{c}}$ 6.78
$[\mathcal{C}_{qu}^{(1)}]_{3333}$ 6.73 $m_{W}$ 1.60 $R_{t}^{365}$ 6.74
$[\mathcal{C}_{qq}^{(3)}]_{1133}$ 5.92 $\Gamma_{Z}$ 3.92 $R_{c}^{240}$ 6.58
$[\mathcal{C}_{lq}^{(3)}]_{3311}$ 4.87 $A^{Z}_{\tau}$ 4.62 $R_{\tau}^{240}$ 6.58
$[\mathcal{C}_{lq}^{(3)}]_{3322}$ 4.79 $A^{Z}_{\tau}$ 4.61 $R_{\tau}^{240}$ 6.54
$[\mathcal{C}_{qq}^{(3)}]_{1122}$ 5.18 $R^{Z}_{\mu}$ 4.38 $R_{e}^{365}$ 6.47
$[\mathcal{C}_{qq}^{(3)}]_{3333}$ 4.79 $R^{Z}_{\mu}$ 4.54 $R_{b}^{240}$ 6.44
$[\mathcal{C}_{uu}]_{3333}$ 6.34 $m_{W}$ 2.05 $R_{t}^{365}$ 6.38
$[\mathcal{C}_{qq}^{(3)}]_{2222}$ 4.44 $R^{Z}_{\mu}$ 4.42 $R_{c}^{240}$ 6.21
$[\mathcal{C}_{qu}^{(1)}]_{2332}$ 0.00 5.87 $\mu_{c\bar{c}}$ 5.87
$[\mathcal{C}_{qq}^{(3)}]_{2233}$ 4.73 $R^{Z}_{\mu}$ 3.81 $R_{e}^{365}$ 5.73
$[\mathcal{C}_{lequ}^{(1)}]_{2133}$ 0.00 5.10 $\mu_{\mu\bar{\mu}}$ 5.10
$[\mathcal{C}_{lequ}^{(1)}]_{1233}$ 0.00 5.10 $\mu_{\mu\bar{\mu}}$ 5.10
$[\mathcal{C}_{ee}]_{2222}$ 4.33 $A^{Z}_{\mu}$ 3.28 $R_{\mu}^{240}$ 5.07
$[\mathcal{C}_{eH}]_{12}$ 0.00 4.94 $\mu_{\mu\bar{\mu}}$ 4.94
$[\mathcal{C}_{eH}]_{21}$ 0.00 4.94 $\mu_{\mu\bar{\mu}}$ 4.94
$[\mathcal{C}_{quqd}^{(8)}]_{3333}$ 0.00 4.88 $\mu_{b\bar{b}}$ 4.88
$[\mathcal{C}_{Hud}]_{33}$ 4.14 $m_{W}$ 2.64 $\mu_{b\bar{b}}$ 4.46
$[\mathcal{C}_{qq}^{(1)}]_{1133}$ 4.45 $\Gamma_{Z}$ 1.04 $R_{t}^{365}$ 4.46
$[\mathcal{C}_{uH}]_{21}$ 0.00 4.45 $\mu_{c\bar{c}}$ 4.45
$[\mathcal{C}_{ee}]_{3333}$ 3.01 $A^{Z}_{\tau}$ 3.28 $R_{\tau}^{240}$ 4.45
$[\mathcal{C}_{qu}^{(1)}]_{1133}$ 4.08 $\Gamma_{Z}$ 1.16 $R_{t}^{365}$ 4.09
$[\mathcal{C}_{eW}]_{33}$ 2.07 $A^{Z}_{e}$ 3.44 $\mu_{\tau\bar{\tau}}$ 4.02
$[\mathcal{C}_{uu}]_{1111}$ 1.17 $R^{Z}_{\mu}$ 3.11 $R_{c}^{240}$ 3.98
$[\mathcal{C}_{dH}]_{32}$ 0.00 3.91 $\mu_{b\bar{b}}$ 3.91
$[\mathcal{C}_{dH}]_{23}$ 0.00 3.86 $\mu_{b\bar{b}}$ 3.86
$[\mathcal{C}_{dH}]_{13}$ 0.00 3.74 $\mu_{b\bar{b}}$ 3.74
$[\mathcal{C}_{qq}^{(1)}]_{2233}$ 3.70 $\Gamma_{Z}$ 1.26 $R_{c}^{240}$ 3.73
$[\mathcal{C}_{eu}]_{2211}$ 3.21 $A^{Z}_{\mu}$ 2.30 $R_{\mu}^{240}$ 3.67
$[\mathcal{C}_{eu}]_{2222}$ 3.21 $A^{Z}_{\mu}$ 2.25 $R_{\mu}^{240}$ 3.64
$[\mathcal{C}_{qu}^{(1)}]_{2233}$ 3.46 $\Gamma_{Z}$ 1.66 $R_{c}^{240}$ 3.58
$[\mathcal{C}_{qu}^{(1)}]_{3311}$ 3.47 $R^{Z}_{\mu}$ 1.60 $R_{t}^{365}$ 3.58
$[\mathcal{C}_{qu}^{(1)}]_{3322}$ 3.45 $R^{Z}_{\mu}$ 1.47 $R_{t}^{365}$ 3.53
$[\mathcal{C}_{dB}]_{33}$ 3.34 $A^{Z}_{e}$ 1.99 $\mu_{b\bar{b}}$ 3.52
[…]

### v2

@@ B.1 Warsaw Basis
Table 4: Single Operator $2\sigma$ confidence intervals in the Warsaw basis exceding $1\text{\,}\mathrm{TeV}$ from On- and Off-Pole observables, as well as their combined projected sensitivity. The Wilson Coefficients are renormalized at $\Lambda=$1\text{\,}\mathrm{TeV}$\$. We also report the most constraining measurement for each coefficient for On- and Off-Pole Observables.
$\mathcal{C}_{\rm{i}}$ $\Lambda^{\rm{eff}}_{\rm{i}}$ on-pole Obs on-pole $\Lambda^{\rm{eff}}_{\rm{i}}$ off-pole Obs off-pole $\Lambda^{\rm{eff}}_{\rm{i}}$ Combined
$\mathcal{C}_{HWB}$ 111.55 $A^{Z}_{e}$ 7.01 $\sigma_{Zh}^{240}$ 111.55
$\mathcal{C}_{Hl}^{(3)}]_{22}$ 67.88 $A^{Z}_{e}$ 26.34 $\sigma_{WW}^{240}$ 68.25
$\mathcal{C}_{Hl}^{(3)}]_{11}$ 65.11 $A^{Z}_{\mu}$ 28.34 $\sigma_{WW}^{240}$ 65.69
$\mathcal{C}_{HD}$ 57.11 $m_{W}$ 4.47 $\sigma_{WW}^{240}$ 57.11
$\mathcal{C}_{ll}]_{1221}$ 52.98 $A^{Z}_{e}$ 37.78 $R_{\mu}^{240}$ 56.13
$\mathcal{C}_{He}]_{11}$ 54.23 $A^{Z}_{e}$ 13.48 $\sigma_{Zh}^{365}$ 54.28
$\mathcal{C}_{He}]_{22}$ 53.58 $A^{Z}_{\mu}$ 1.41 $R_{\mu}^{163}$ 53.58
$\mathcal{C}_{Hl}^{(1)}]_{22}$ 51.18 $A^{Z}_{\mu}$ 0.94 $R_{\mu}^{163}$ 51.18
$\mathcal{C}_{Hl}^{(1)}]_{11}$ 50.85 $A^{Z}_{e}$ 14.97 $\sigma_{Zh}^{365}$ 50.94
$\mathcal{C}_{lq}^{(1)}]_{1133}$ 12.52 $A^{Z}_{e}$ 48.63 $R_{b}^{240}$ 48.68
$\mathcal{C}_{lq}^{(3)}]_{1133}$ 17.12 $A^{Z}_{\mu}$ 47.35 $R_{b}^{240}$ 47.55
$\mathcal{C}_{lq}^{(3)}]_{1111}$ 6.84 $A^{Z}_{e}$ 47.49 $R_{c}^{240}$ 47.50
$\mathcal{C}_{lq}^{(1)}]_{1122}$ 2.12 $A^{Z}_{e}$ 46.94 $R_{c}^{240}$ 46.94
$\mathcal{C}_{lq}^{(3)}]_{1122}$ 6.81 $A^{Z}_{e}$ 45.57 $R_{c}^{240}$ 45.57
$\mathcal{C}_{He}]_{33}$ 37.27 $A^{Z}_{\tau}$ 1.41 $R_{\tau}^{163}$ 37.27
$\mathcal{C}_{Hl}^{(3)}]_{33}$ 36.59 $A^{Z}_{\tau}$ 2.21 $R_{\tau}^{163}$ 36.59
$\mathcal{C}_{eu}]_{1122}$ 3.26 $A^{Z}_{e}$ 36.36 $R_{c}^{240}$ 36.36
$\mathcal{C}_{Hl}^{(1)}]_{33}$ 35.72 $A^{Z}_{\tau}$ 0.94 $R_{\tau}^{163}$ 35.72
$\mathcal{C}_{ll}]_{1122}$ 11.94 $A^{Z}_{e}$ 35.29 $R_{\mu}^{240}$ 35.41
$\mathcal{C}_{ll}]_{1133}$ 2.30 $A^{Z}_{e}$ 35.29 $R_{\tau}^{240}$ 35.29
$\mathcal{C}_{ll}]_{1331}$ 2.25 $A^{Z}_{e}$ 35.19 $R_{\tau}^{240}$ 35.19
$\mathcal{C}_{ee}]_{1122}$ 2.67 $A^{Z}_{e}$ 33.10 $R_{\mu}^{240}$ 33.10
$\mathcal{C}_{ee}]_{1133}$ 2.38 $A^{Z}_{e}$ 33.10 $R_{\tau}^{240}$ 33.10
$\mathcal{C}_{Hq}^{(3)}]_{11}$ 31.01 $\Gamma_{Z}$ 1.90 $\sigma_{Zh}^{240}$ 31.01
$\mathcal{C}_{Hq}^{(1)}]_{33}$ 30.78 $m_{W}$ 1.87 $\sigma_{WW}^{240}$ 30.78
$\mathcal{C}_{ed}]_{1133}$ 2.27 $A^{Z}_{e}$ 30.65 $R_{b}^{240}$ 30.65
$\mathcal{C}_{eu}]_{1111}$ 3.26 $A^{Z}_{e}$ 29.60 $R_{e}^{365}$ 29.61
$\mathcal{C}_{qe}]_{3311}$ 13.20 $A^{Z}_{e}$ 26.50 $R_{b}^{240}$ 26.90
$\mathcal{C}_{Hu}]_{33}$ 26.04 $m_{W}$ 1.94 $\sigma_{WW}^{240}$ 26.04
$\mathcal{C}_{Hq}^{(3)}]_{22}$ 25.60 $R^{Z}_{\mu}$ 1.86 $\sigma_{Zh}^{240}$ 25.60
$\mathcal{C}_{Hq}^{(3)}]_{33}$ 23.70 $R^{Z}_{\mu}$ 1.70 $R_{b}^{240}$ 23.70
$\mathcal{C}_{lu}]_{1122}$ 3.04 $A^{Z}_{e}$ 23.66 $R_{c}^{240}$ 23.66
$\mathcal{C}_{eu}]_{1133}$ 12.39 $A^{Z}_{e}$ 21.90 $R_{t}^{365}$ 22.44
$\mathcal{C}_{uW}]_{33}$ 22.29 $A^{Z}_{e}$ 1.15 $\sigma_{Zh}^{365}$ 22.29
$\mathcal{C}_{uB}]_{33}$ 22.04 $A^{Z}_{e}$ 0.99 $\sigma_{WW}^{240}$ 22.04
$\mathcal{C}_{le}]_{2211}$ 2.65 $A^{Z}_{e}$ 21.54 $R_{\mu}^{240}$ 21.54
$\mathcal{C}_{le}]_{3311}$ 2.39 $A^{Z}_{e}$ 21.54 $R_{\tau}^{240}$ 21.54
$\mathcal{C}_{le}]_{1122}$ 2.62 $A^{Z}_{\mu}$ 21.52 $R_{\mu}^{240}$ 21.52
$\mathcal{C}_{le}]_{1133}$ 2.29 $A^{Z}_{e}$ 21.52 $R_{\tau}^{240}$ 21.52
$\mathcal{C}_{lu}]_{1133}$ 11.78 $A^{Z}_{e}$ 20.95 $R_{t}^{365}$ 21.46
$\mathcal{C}_{ed}]_{1122}$ 2.28 $A^{Z}_{e}$ 20.86 $R_{e}^{365}$ 20.86
$\mathcal{C}_{qe}]_{1111}$ 2.30 $A^{Z}_{e}$ 20.82 $R_{e}^{365}$ 20.82
$\mathcal{C}_{ed}]_{1111}$ 2.29 $A^{Z}_{e}$ 20.80 $R_{e}^{365}$ 20.80
$\mathcal{C}_{ld}]_{1133}$ 2.14 $A^{Z}_{e}$ 20.31 $R_{b}^{240}$ 20.31
$\mathcal{C}_{ll}]_{1111}$ 1.06 $\Gamma_{Z}$ 19.42 $R_{e}^{365}$ 19.42
$\mathcal{C}_{lu}]_{1111}$ 3.04 $A^{Z}_{e}$ 19.41 $R_{e}^{365}$ 19.42
$\mathcal{C}_{ee}]_{1111}$ 4.39 $A^{Z}_{e}$ 19.23 $R_{e}^{365}$ 19.25
$\mathcal{C}_{qe}]_{2211}$ 2.38 $A^{Z}_{e}$ 19.20 $R_{e}^{365}$ 19.21
$\mathcal{C}_{lq}^{(3)}]_{2233}$ 17.55 $A^{Z}_{e}$ 7.04 $\sigma_{WW}^{161}$ 17.67
$\mathcal{C}_{Hq}^{(1)}]_{11}$ 17.64 $\Gamma_{Z}$ 0.68 $\sigma_{WW}^{240}$ 17.64
$\mathcal{C}_{le}]_{1111}$ 3.12 $A^{Z}_{e}$ 15.94 $R_{e}^{365}$ 15.94
$\mathcal{C}_{Hq}^{(1)}]_{22}$ 15.06 $\Gamma_{Z}$ 0.94 $R_{c}^{240}$ 15.06
$\mathcal{C}_{lq}^{(1)}]_{1111}$ 2.04 $A^{Z}_{e}$ 14.58 $R_{e}^{365}$ 14.58
$\mathcal{C}_{ld}]_{1122}$ 2.16 $A^{Z}_{e}$ 13.81 $R_{e}^{365}$ 13.82
$\mathcal{C}_{ld}]_{1111}$ 2.16 $A^{Z}_{e}$ 13.77 $R_{e}^{365}$ 13.77
$\mathcal{C}_{Hu}]_{22}$ 13.17 $R^{Z}_{\mu}$ 1.14 $R_{c}^{240}$ 13.17
$\mathcal{C}_{Hu}]_{11}$ 13.10 $R^{Z}_{\mu}$ 0.95 $\sigma_{WW}^{240}$ 13.10
$\mathcal{C}_{qe}]_{3322}$ 13.05 $A^{Z}_{\mu}$ 2.04 $R_{\mu}^{163}$ 13.06
$\mathcal{C}_{lq}^{(1)}]_{2233}$ 12.59 $A^{Z}_{\mu}$ 1.39 $R_{\mu}^{163}$ 12.59
$\mathcal{C}_{eu}]_{2233}$ 12.26 $A^{Z}_{\mu}$ 2.83 $R_{\mu}^{163}$ 12.27
$\mathcal{C}_{lu}]_{2233}$ 11.81 $A^{Z}_{\mu}$ 1.87 $R_{\mu}^{163}$ 11.82
$\mathcal{C}_{HW}$ 9.29 $A^{Z}_{e}$ 9.21 $\sigma_{Zh}^{240}$ 11.00
$\mathcal{C}_{Hd}]_{11}$ 10.85 $\Gamma_{Z}$ 0.67 $\sigma_{WW}^{240}$ 10.85
$\mathcal{C}_{Hd}]_{33}$ 9.64 $R^{Z}_{\mu}$ 0.94 $R_{b}^{240}$ 9.64
$\mathcal{C}_{Hq}^{(1)}]_{23}$ 9.32 $m_{W}$ 0.55 $\sigma_{WW}^{240}$ 9.32
$\mathcal{C}_{HB}$ 9.16 $A^{Z}_{e}$ 4.71 $\sigma_{Zh}^{240}$ 9.31
$\mathcal{C}_{qe}]_{3333}$ 9.08 $A^{Z}_{\tau}$ 2.04 $R_{\tau}^{163}$ 9.09
$\mathcal{C}_{W}$ 8.63 $A^{Z}_{e}$ 5.52 $\sigma_{WW}^{240}$ 8.97
$\mathcal{C}_{lq}^{(1)}]_{3333}$ 8.82 $A^{Z}_{\tau}$ 1.38 $R_{\tau}^{163}$ 8.82
$\mathcal{C}_{qq}^{(1)}]_{3333}$ 8.69 $R^{Z}_{\mu}$ 2.00 $R_{t}^{365}$ 8.70
$\mathcal{C}_{eu}]_{3333}$ 8.53 $A^{Z}_{\tau}$ 2.83 $R_{\tau}^{163}$ 8.56
$\mathcal{C}_{Hd}]_{22}$ 8.48 $R^{Z}_{\mu}$ 0.67 $\sigma_{WW}^{240}$ 8.48
$\mathcal{C}_{lu}]_{3333}$ 8.25 $A^{Z}_{\tau}$ 1.87 $R_{\tau}^{163}$ 8.25
$\mathcal{C}_{lq}^{(3)}]_{3333}$ 7.63 $A^{Z}_{\tau}$ 5.50 $R_{\tau}^{163}$ 8.10
$\mathcal{C}_{lq}^{(3)}]_{2211}$ 6.82 $A^{Z}_{\mu}$ 5.65 $R_{\mu}^{163}$ 7.51
$\mathcal{C}_{lq}^{(3)}]_{2222}$ 6.79 $A^{Z}_{\mu}$ 5.65 $R_{\mu}^{163}$ 7.49
$\mathcal{C}_{H\square}$ 6.84 $m_{W}$ 3.96 $\sigma_{Zh}^{240}$ 7.02
$\mathcal{C}_{qu}^{(1)}]_{3333}$ 6.72 $m_{W}$ 1.60 $R_{t}^{365}$ 6.73
$\mathcal{C}_{qq}^{(3)}]_{1133}$ 5.91 $\Gamma_{Z}$ 4.51 $R_{c}^{240}$ 6.36
$\mathcal{C}_{qq}^{(3)}]_{1111}$ 5.42 $\Gamma_{Z}$ 5.26 $R_{c}^{240}$ 6.36
$\mathcal{C}_{uu}]_{3333}$ 6.34 $m_{W}$ 2.01 $R_{t}^{365}$ 6.35
$\mathcal{C}_{lq}^{(3)}]_{3311}$ 4.87 $A^{Z}_{\tau}$ 5.61 $R_{\tau}^{163}$ 6.27
$\mathcal{C}_{lq}^{(3)}]_{3322}$ 4.77 $A^{Z}_{\tau}$ 5.60 $R_{\tau}^{163}$ 6.23
$\mathcal{C}_{qq}^{(3)}]_{1122}$ 5.17 $R^{Z}_{\mu}$ 5.05 $R_{\tau}^{163}$ 6.08
$\mathcal{C}_{qq}^{(3)}]_{3333}$ 4.79 $R^{Z}_{\mu}$ 5.17 $R_{b}^{240}$ 5.94
$\mathcal{C}_{qq}^{(3)}]_{2222}$ 4.42 $R^{Z}_{\mu}$ 5.10 $R_{c}^{163}$ 5.70
$\mathcal{C}_{qq}^{(3)}]_{2233}$ 4.72 $R^{Z}_{\mu}$ 4.40 $R_{\tau}^{163}$ 5.43
$\mathcal{C}_{uB}]_{23}$ 5.26 $A^{Z}_{e}$ 0.00 5.26
$\mathcal{C}_{uW}]_{23}$ 5.11 $A^{Z}_{e}$ 0.00 5.11
$\mathcal{C}_{ee}]_{2222}$ 4.33 $A^{Z}_{\mu}$ 3.96 $R_{\mu}^{163}$ 4.95
$\mathcal{C}_{qq}^{(1)}]_{1133}$ 4.44 $\Gamma_{Z}$ 1.12 $R_{b}^{240}$ 4.45
$\mathcal{C}_{ee}]_{3333}$ 3.01 $A^{Z}_{\tau}$ 3.96 $R_{\tau}^{163}$ 4.26
$\mathcal{C}_{Hq}^{(1)}]_{13}$ 4.20 $m_{W}$ 0.00 4.20
$\mathcal{C}_{Hud}]_{33}$ 4.14 $m_{W}$ 0.00 4.14
$\mathcal{C}_{qu}^{(1)}]_{1133}$ 4.08 $\Gamma_{Z}$ 1.30 $R_{t}^{365}$ 4.09
$\mathcal{C}_{qq}^{(1)}]_{2233}$ 3.69 $\Gamma_{Z}$ 1.39 $R_{c}^{240}$ 3.71
$\mathcal{C}_{eu}]_{2222}$ 3.21 $A^{Z}_{\mu}$ 2.77 $R_{\mu}^{163}$ 3.58
$\mathcal{C}_{eu}]_{2211}$ 3.21 $A^{Z}_{\mu}$ 2.72 $R_{\mu}^{163}$ 3.56
$\mathcal{C}_{uu}]_{2222}$ 1.17 $R^{Z}_{\mu}$ 3.55 $R_{c}^{240}$ 3.56
$\mathcal{C}_{qu}^{(1)}]_{2233}$ 3.46 $\Gamma_{Z}$ 1.87 $R_{c}^{240}$ 3.53
$\mathcal{C}_{qu}^{(1)}]_{3322}$ 3.47 $R^{Z}_{\mu}$ 1.76 $R_{c}^{240}$ 3.52
[…]